The length of a rectangle is four inches longer than twice its width. If both dimensions were increased by the length equal to its width, the perimeter would be three inches longer than 3/2 times its original perimeter. Find the dimensions of the original rectangle and its perimeter.
I tried
L=4+2w
LLW+WWL=3+(3/2+P)
I'm absolutely puzzled. please help me out!!

Answer by richwmiller(17219)

"If both dimensions were increased by the length equal to its width"
Maybe it means all dimensions were increased by the original width
10w+8=3+3/2(6w+8)=new perimeter
w=7 old width
L=4+2*7=18 old length
14+36=50 old perimeter
new dimensions
nw=7+7=14
nl=18+7=25
28+50=78 new perimeter